OpenStudy (abbycross167):

Can someone please help me with this question? I will fan and give a medal... Simplify (3a-4)-(7a-4) a. -4a - 8 b. -4a c. -4a + 8 d. 10a

OpenStudy (welshfella):

the tricjiest part is to distribute the negative over the (7a - 4) its the same as -1(7a - 4) which equals -1*7a -1*-4

OpenStudy (welshfella):

so we have 3a - 4 -1*7a -1*-4 can you continue?

OpenStudy (abbycross167):

1*7=7 1*-4=-4 3a-4-7a-4 3a-7a-4-4 so a.-4a - 8?

OpenStudy (welshfella):

no -1* 7a = -7a ( - * + = -) -1*-4 = + 4 ( when 2 negatives are multiplied the answer is a positive)

OpenStudy (abbycross167):

ooh so would it be c.-4a + 8?

OpenStudy (welshfella):

3a - 4 - 7a + 4 - bring like terms together = 3a - 7a - 4 + 4

OpenStudy (welshfella):

3a - 71 = -4a that is correct but its not + 8 . Can you work out - 4 + 4?

OpenStudy (welshfella):

* 3a - 7a = -4a

OpenStudy (welshfella):

-4 + 4 = 4 - 4 = ?

OpenStudy (abbycross167):

0

OpenStudy (welshfella):

correct so which is the correct choice?

OpenStudy (abbycross167):

b. -4a

OpenStudy (welshfella):

we needn't write 0 yes -4a

OpenStudy (abbycross167):

oh ok, thank you very much!! you're very helpful!

OpenStudy (welshfella):

yw